# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v1

@@ 7.1 Classification and Regression on Incomplete data
The best-known strategy for classification and regression on incomplete data is the EM algorithm (Dempster et al., 1977), which was introduced in Section 5.4. There have been more methods for classification and regression on incomplete data. García-Laencina et al. (2009a) reviewed different ways to handle missing data in classification tasks, grouping methods into categories like deleting incomplete cases, imputing (estimating) missing values, and using models that can work with incomplete data directly. The following discussion organizes the methods of classification and regression on incomplete data by model type, progressing from linear to non-linear.
@@
Williams et al. (2005) presented a logistic regression-based classification method for datasets with missing features, avoiding imputation by analytically integrating over the missing values using a Gaussian Mixture Model (GMM). Goldberg et al. (2010) concatenated the feature matrix and label matrix, treated the unknown labels as a part of missing values, and used low-rank matrix completion to predict the unknown labels directly, where the classifier is linear. Hazan et al. (2015) proposed an efficient, theoretically grounded algorithm that learns a linear classifier directly from incomplete data without requiring matrix completion. The method uses a kernel-based approach to embed observed data into a high-dimensional space, enabling competitive classification performance with the best possible linear classifier trained on full data, even when most features are missing.
@@
García-Laencina et al. (2009b) proposed a kNN method with mutual information for simultaneous classification and missing data imputation. Sell et al. (2024) introduced a nonparametric framework for classification problems in the presence of missing data and derived the minimax rate for the excess risk of classification. It also proposed a Hard-thresholding Anova Missing data (HAM) classifier based on kNN and a thresholding step.
@@
Twala et al. (2008) proposed a method called MIA for handling missing data in decision trees by treating missingness as a separate, informative group during the splitting process. The authors demonstrated that MIA performs competitively with more complex methods like the EM-based method, while offering advantages in simplicity and computational efficiency. Hapfelmeier et al. (2014) introduced a new variable importance measure that is applicable to data with or without missing values, thereby enabling the use of random forests on incomplete datasets.
@@
Smola et al. (2005) introduced a framework for handling missing data in kernel methods like Support Vector Machines (SVM) (Boser et al., 1992) and Gaussian Processes (Williams and Rasmussen, 1995) by modeling them as estimation problems within exponential families. The key idea is to treat estimation with missing variables as a problem of computing marginal distributions, which leads to non-convex optimization problems. Chechik et al. (2006) developed a max-margin classifier that works directly with incomplete data, avoiding the need for a separate imputation step. Their method uses a geometry-based approach with different optimization strategies for separable and non-separable cases. Fan et al. (2020a) proposed to use polynomial matrix completion to do nonlinear classification on incomplete data, showing improved performance compared to the method of (Goldberg et al., 2010), SVM, and LRMC+SVM.
@@
Although deep learning-based imputation methods (Section 5.6) are often adaptable to joint imputation and classification, some research focuses specifically on classification without imputation or prioritizes the supervised learning objective. For instance, Ghorbani and Zou (2018) proposed an embedding approach to learn representations for missingness, where the embedding is learned at the same time as the networks learn to make predictions. The approach bypasses the need to impute the missing attributes. For incomplete image classification, Danel et al. (2020) proposed to convert images with missing values to graphs, and then used GCN to do graph classification, which avoids imputation. Chang et al. (2022) presented GapNet, a neural network method that can be applied to highly incomplete medical datasets. Kim et al. (2023) developed a probabilistic framework for classifying multivariate time series data with missing values by combining a deep generative model for imputation with a classifier. To prevent the model from taking the easy way out and ignoring the imputed values, the authors introduce a novel regularization technique called ‘obsdropout’, which randomly drops observed values during training to force the classifier to rely more on the generated imputations.
@@ 7.2 Clustering on Incomplete Data
Clustering is an unsupervised learning task. Therefore, handling missing data in clustering is more challenging than in classification and regression. In the past decades, many clustering algorithms have been proposed. Typical clustering algorithms include k-means (Ball and Hall, 1965; Ikotun et al., 2023), DBSCAN (Ester et al., 1996), Gaussian mixture models, spectral clustering (Shi and Malik, 2000; Ng et al., 2001), subspace clustering (Parsons et al., 2004; Elhamifar and Vidal, 2013), and deep learning-based clustering (Xie et al., 2016a; Ji et al., 2017; Cai et al., 2022b; Ren et al., 2024b). When the data are incomplete, we can use some missing data imputation methods to fill the missing values and then perform these clustering algorithms. However, such a two-stage approach may not work well if the imputation algorithm does not align with the cluster structure in the data. There have been a few efforts in designing more effective approaches for clustering on incomplete data.
@@
Sarkar and Leong (2001) proposed an algorithm for fuzzy k-means on incomplete data, where the missing values are repaired incrementally in each iteration. Wagstaff (2004) presented a method for encoding partially observed features as a set of supplemental soft constraints and introduced an algorithm of k-means with soft constraints, which incorporates constraints into the clustering process, thereby eliminating the need of imputation. Chi et al. (2016) developed an algorithm called k-POD for k-means with missing data. k-POD minimizes the sum of the squared differences between the data and the centers over the observed entries only, remaining unhindered by the need for an accurate imputations. Wang et al. (2019) proposed a variant of k-means that unifies the clustering and imputation into one single objective, where the missing values are optimized dynamically. The method outperformed k-means with kNN, EM, and mean imputations on a few datasets. Armah and Melnykov (2024) extended the idea to k-means with the Mahalanobis distance. There are also studies on kernel k-means for incomplete data such as (Liu et al., 2020) and Gaussian mixture models with missing data (Zhang et al., 2021a).
@@
Spectral clustering first constructs the similarity matrix and then performs k-means on the representations given by the eigenvalue decomposition of the Laplacian matrix (Ng et al., 2001). A simple strategy to handling missing data is computing the similarity between two data points using only the observed values, especially when the Gaussian kernel is used. However, this may not work well when the missing rate is high. Løkse et al. (2017) developed a robust kernel function that is specifically designed to handle incomplete data and applied it to spectral clustering. Chen et al. (2024a) presented a spectral embedding fusion method for incomplete multi-view clustering. Wen et al. (2022) wrote a survey paper of incomplete multi-view clustering, which included a few extension methods of spectral clustering for incomplete data. Yu et al. (2023) proposed an imputation-free framework with two approaches to improving spectral clustering on incomplete data. The first approach is based on a kernel correction method, that finds the best positive semi-definite approximation for the kernel matrix estimated on incomplete data. The second one is an extension of the first one, by using kernel self-expressive learning.
@@
Most subspace clustering methods can be organized into two categories. The first category is based on learning the subspaces directly (Bradley and Mangasarian, 2000; Fan, 2021), while the second category is based on self-expressive models (Elhamifar and Vidal, 2013) followed by spectral clustering. Yang et al. (2015) developed two methods for subspace clustering on incomplete data. The first one generalizes the sparse subspace clustering algorithm so that it can obtain a sparse representation of the data using only the observed entries. The second one estimates a suitable kernel matrix by assuming a random model for the missing entries and obtains the sparse representation from this kernel. Pimentel-Alarcón et al. (2016) also proposed two methods for subspace clustering on incomplete data: (a) group-sparse subspace clustering (GSSC), which is based on group-sparsity and alternating minimization, and (b) mixture subspace clustering (MSC), which models each data point as a convex combination of its projections onto all subspaces in the union. Pimentel-Alarcon and Nowak (2016) analyzed the information-theoretic requirements of subspace clustering with missing data. Fan and Chow (2017c) proposed to alternately compute the matrix of sparse representation coefficients and recover the missing entries of a data matrix. In particular, the algorithm recovers missing entries by minimizing the representation coefficients, representation errors, and matrix rank. Tsakiris and Vidal (2018) provided theoretical guarantees for sparse subspace clustering (Elhamifar and Vidal, 2013) with incomplete data, showing that projecting the zero-filled data onto the observation pattern of the point being expressed can lead to substantial improvement in performance. Lane et al. (2019) provided a comprehensive comparison of methods for subspace clustering with missing data. Wang et al. (2020a) proposed a method for incomplete cross-modal subspace clustering. Liu et al. (2021) developed an algorithm for self-representation subspace clustering on incomplete multi-view data. Soni et al. (2025) proposed to dynamically determine a set of candidate subspaces and optimally assign points to selected subspaces. The empirical study showed that the proposed approach can achieve high clustering accuracy even when the data are high rank, the percentage of missing data is high, or the subspaces are similar.
@@
de Jong et al. (2019) developed a variational deep embedding model for clustering of multivariate clinical patient trajectories with missing values. Wen et al. (2020) proposed a deep learning model for incomplete multi-view clustering. Xu et al. (2024b) provided a deep variational incomplete multi-view clustering algorithm that is able to explore shared clustering structures of incomplete data from different views. More work on deep clustering with missing data can be found in (Zhao et al., 2018; Wen et al., 2021; Xu et al., 2022; Lin et al., 2022; Tang and Liu, 2022).
@@ 7.3 Anomaly/Outlier/Novelty Detection on Incomplete Data
Anomaly detection (Chandola et al., 2009), sometimes known as outlier detection (Boukerche et al., 2020), novelty detection (Pimentel et al., 2014), or even fault detection (in the manufacturing industry) (Venkatasubramanian et al., 2003), is usually an unsupervised learning task that aims to identify the samples deviating from common patterns. Typical algorithms include kernel density estimation, one-class SVM (Schölkopf et al., 1999), Isolation Forest (Liu et al., 2008), Deep SVDD (Ruff et al., 2018), etc. Note that most of the dimensionality reduction methods (Van Der Maaten et al., 2009), such as PCA (Abdi and Williams, 2010), KPCA (Schölkopf et al., 1998), and autoencoder (Hinton and Salakhutdinov, 2006), can be used for anomaly detection or outlier detection. More papers about anomaly detection, especially those based on deep learning, can be found in (Pang et al., 2021).
Zemicheal and Dietterich (2019) evaluated the detection performance of a few anomaly detection methods combined with different data imputation techniques. The experimental results showed that implementations contribute positively to the detection performance of unsupervised anomaly detection methods on incomplete data. Fan et al. (2022) investigated statistical process monitoring in the presence of missing data and introduced a fast incremental nonlinear matrix completion technique for online and sequential imputation. Xu and Xie (2021) developed ECAD, a method tailored for incomplete spatio-temporal data like traffic flow sequences, while Ma et al. (2021) utilized Probabilistic Principal Component Analysis (PPCA) to recover missing values and identify anomalies. Sarda et al. (2023) conducted a comparative analysis of existing unsupervised anomaly detection techniques applied to GAN-imputed datasets. These approaches follow a two-stage framework, in which the imputation models are trained exclusively on normal data or datasets with very few unlabeled outliers. Consequently, when applied to abnormal data during inference, the imputation model tends to reconstruct missing values based on learned normal patterns, thereby making anomalies appear more normal and reducing detection accuracy.
Xiao and Fan (2024) proposed an end-to-end method called ImAD for unsupervised anomaly detection in the presence of missing values. It integrates data imputation with anomaly detection into a unified optimization objective and automatically generates pseudo-abnormal samples to alleviate the imputation bias. They also proved the effectiveness of ImAD theoretically. Zhang et al. (2025f) proposed an anomaly detection method for incomplete time series that flags unusual patterns by checking how consistently a two-step, patch-based imputation network fills the same series under different random masks.
@@ 7.4 Other Tasks on Incomplete Data
(Sanguinetti and Lawrence, 2006; Nguyen and Torre, 2008; Hu et al., 2010)
(De Bodt et al., 2018; Gilbert and Sonthalia, 2018; Ling et al., 2021; Yan et al., 2024)
@@
(Xie et al., 2012b; Londschien et al., 2021; Zhao et al., 2022; Follain et al., 2022)
(Liu and Safikhani, 2025; Xu et al., 2025; Enikeeva and Klopp, 2025)
@@
(Kallus et al., 2018; Yang et al., 2019; Tu et al., 2019; Nguyen and Gouno, 2020)
(Wang et al., 2020b; Huang et al., 2020b; Hillis et al., 2021)
(Gao et al., 2022; Levis et al., 2025; Landsiedel et al., 2025)
@@ 8.1 Effective Parameter Estimation
Dempster et al. (1977) proved that the EM algorithm for handling missing data converges to a local maximum point of the likelihood function. Although this does not directly guarantee accurate missing data imputation, it guarantees the effectiveness of the EM algorithm in estimating the parameters of many models such as mixture models and robust regression. Chandrasekher et al. (2020) studied the problem of single imputation for high-dimensional sparse linear regression and obtained optimal consistency rates of both the LASSO (Ranstam and Cook, 2018) and the square-root LASSO without modification. Bertsimas et al. (2024) studied the performance of impute-then-regress pipelines by contrasting theoretical and empirical evidence. It established the asymptotic consistency of such pipelines for a broad family of imputation methods and showed that mean-impute is asymptotically optimal. Verchand and Montanari (2024) presented an asymptotically exact characterization of the risk of ridge-regularized logistic regression when the observed covariates stem from a Gaussian error-in-variables model.

### v2

@@ 7.1 Classification and Regression on Incomplete data
The best-known strategy for classification and regression on incomplete data is the EM algorithm (dempster1977maximum), which was introduced in Section 5.4. There have been more methods for classification and regression on incomplete data. EM reviewed different ways to handle missing data in classification tasks, grouping methods into categories like deleting incomplete cases, imputing (estimating) missing values, and using models that can work with incomplete data directly. The following discussion organizes the methods of classification and regression on incomplete data by model type, progressing from linear to non-linear.
@@
williams2005incomplete presented a logistic regression-based classification method for datasets with missing features, avoiding imputation by analytically integrating over the missing values using a Gaussian Mixture Model (GMM). NIPS2010_3932 concatenated the feature matrix and label matrix, treated the unknown labels as a part of missing values, and used low-rank matrix completion to predict the unknown labels directly, where the classifier is linear. hazan2015classification proposed an efficient, theoretically grounded algorithm that learns a linear classifier directly from incomplete data without requiring matrix completion. The method uses a kernel-based approach to embed observed data into a high-dimensional space, enabling competitive classification performance with the best possible linear classifier trained on full data, even when most features are missing.
@@
garcia2009k proposed a kNN method with mutual information for simultaneous classification and missing data imputation. sell2024nonparametric introduced a nonparametric framework for classification problems in the presence of missing data and derived the minimax rate for the excess risk of classification. It also proposed a Hard-thresholding Anova Missing data (HAM) classifier based on kNN and a thresholding step.
@@
TWALA2008950 proposed a method called MIA for handling missing data in decision trees by treating missingness as a separate, informative group during the splitting process. The authors demonstrated that MIA performs competitively with more complex methods like the EM-based method, while offering advantages in simplicity and computational efficiency. hapfelmeier2014new introduced a new variable importance measure that is applicable to data with or without missing values, thereby enabling the use of random forests on incomplete datasets.
@@
smola2005kernel introduced a framework for handling missing data in kernel methods like Support Vector Machines (SVM) (boser1992training) and Gaussian Processes (williams1995gaussian) by modeling them as estimation problems within exponential families. The key idea is to treat estimation with missing variables as a problem of computing marginal distributions, which leads to non-convex optimization problems. chechik2006max developed a max-margin classifier that works directly with incomplete data, avoiding the need for a separate imputation step. Their method uses a geometry-based approach with different optimization strategies for separable and non-separable cases. fan2020polynomial proposed to use polynomial matrix completion to do nonlinear classification on incomplete data, showing improved performance compared to the method of (NIPS2010_3932), SVM, and LRMC+SVM.
@@
Although deep learning-based imputation methods (Section 5.6) are often adaptable to joint imputation and classification, some research focuses specifically on classification without imputation or prioritizes the supervised learning objective. For instance, ghorbani2018embedding proposed an embedding approach to learn representations for missingness in parallel with the network’s prediction task. The approach bypasses the need to impute the missing attributes. For incomplete image classification, danel2020processing proposed to convert images with missing values to graphs, and then use GCN to do graph classification, which avoids imputation. chang2022neural presented GapNet, a neural network method that can be applied to highly incomplete medical datasets. kim2023probabilistic developed a probabilistic framework for classifying multivariate time series data with missing values by combining a deep generative model for imputation with a classifier. To prevent the model from taking the easy way out and ignoring the imputed values, the authors introduce a novel regularization technique called ‘obsdropout’, which randomly drops observed values during training to force the classifier to rely more on the generated imputations.
@@ 7.2 Clustering on Incomplete Data
Clustering is an unsupervised learning task. Therefore, handling missing data in clustering is more challenging than in classification and regression. In recent decades, many clustering algorithms have been proposed. Typical clustering algorithms include k-means (ball1965isodata; IKOTUN2023178), DBSCAN (ester1996density), Gaussian mixture models, spectral clustering (shi2000normalized; ng2001spectral), subspace clustering (parsons2004subspace; elhamifar2013sparse), and deep learning-based clustering (xie2016unsupervised; ji2017deep; cai2022efficient; ren2024deep). When the data are incomplete, we can use some missing data imputation methods to fill the missing values and then perform these clustering algorithms. However, such a two-stage approach may not work well if the imputation algorithm does not align with the cluster structure in the data. There have been a few efforts in designing more effective approaches for clustering on incomplete data.
@@
sarkar2001fuzzy proposed an algorithm for fuzzy k-means on incomplete data, where the missing values are repaired incrementally in each iteration. wagstaff2004clustering presented a method for encoding partially observed features as a set of supplemental soft constraints and introduced an algorithm of k-means with soft constraints, which incorporates constraints into the clustering process, thereby eliminating the need for imputation. chi2016k developed an algorithm called k-POD for k-means with missing data. k-POD minimizes the sum of the squared differences between the data and the centers over the observed entries only, remaining unhindered by the need for accurate imputations. 8693952 proposed a variant of k-means that unifies clustering and imputation into one single objective, where missing values are optimized dynamically. The method outperformed k-means with kNN, EM, and mean imputations on a few datasets. armah2024k extended the idea to k-means with the Mahalanobis distance. There are also studies on kernel k-means for incomplete data such as (liu2020efficient) and Gaussian mixture models with missing data (zhang2021gaussian).
@@
Spectral clustering first constructs the similarity matrix and then performs k-means on the representations given by the eigenvalue decomposition of the Laplacian matrix (ng2001spectral). A simple strategy to handling missing data is computing the similarity between two data points using only the observed values, especially when the Gaussian kernel is used. However, this may not work well when the missing rate is high. lokse2017spectral developed a robust kernel function that is specifically designed to handle incomplete data and applied it to spectral clustering. 10586869 presented a spectral embedding fusion method for incomplete multi-view clustering. wen2022survey wrote a survey paper of incomplete multi-view clustering, which included a few extension methods of spectral clustering for incomplete data. NEURIPS2023_e5aa7171 proposed an imputation-free framework with two approaches to improving spectral clustering on incomplete data. The first approach is based on a kernel correction method that finds the best positive semi-definite approximation for the kernel matrix estimated on incomplete data. The second one is an extension of the first one, by using kernel self-expressive learning.
@@
Most subspace clustering methods can be organized into two categories. The first category is based on learning the subspaces directly (bradley2000k; fan2021large), while the second category is based on self-expressive models (elhamifar2013sparse) followed by spectral clustering. yang2015sparse developed two methods for subspace clustering on incomplete data. The first one generalizes the sparse subspace clustering algorithm so that it can obtain a sparse representation of the data using only the observed entries. The second one estimates a suitable kernel matrix by assuming a random model for the missing entries and obtains the sparse representation from this kernel. pimentel2016group also proposed two methods for subspace clustering on incomplete data: (a) group-sparse subspace clustering (GSSC), which is based on group-sparsity and alternating minimization, and (b) mixture subspace clustering (MSC), which models each data point as a convex combination of its projections onto all subspaces in the union. pimentel2016information analyzed the information-theoretic requirements of subspace clustering with missing data. fan2017sparse proposed to alternately compute the matrix of sparse representation coefficients and recover the missing entries of a data matrix. In particular, the algorithm recovers missing entries by minimizing the representation coefficients, representation errors, and matrix rank. tsakiris2018theoretical provided theoretical guarantees for sparse subspace clustering (elhamifar2013sparse) with incomplete data, showing that projecting the zero-filled data onto the observation pattern of the point being expressed can lead to substantial improvement in performance. lane2019classifying provided a comprehensive comparison of methods for subspace clustering with missing data. wang2020icmsc proposed a method for incomplete cross-modal subspace clustering. liu2021self developed an algorithm for self-representation subspace clustering on incomplete multi-view data. soni2025integer proposed to dynamically determine a set of candidate subspaces and optimally assign points to selected subspaces. The empirical study showed that the proposed approach can achieve high clustering accuracy even when the data are high rank, the percentage of missing data is high, or the subspaces are similar.
@@
de2019deep developed a variational deep embedding model for clustering of multivariate clinical patient trajectories with missing values. wen2020dimc proposed a deep learning model for incomplete multi-view clustering. xu2024deep provided a deep variational incomplete multi-view clustering algorithm that is able to explore shared clustering structures of incomplete data from different views. More work on deep clustering with missing data can be found in (ZHAO20181053; wen2021structural; xu2022deep; lin2022incomplete; tang2022deep).
@@ 7.3 Anomaly/Outlier/Novelty Detection on Incomplete Data
Anomaly detection (chandola2009anomaly), sometimes known as outlier detection (boukerche2020outlier), novelty detection (pimentel2014review), or even fault detection (in the manufacturing industry) (venkatasubramanian2003review), is usually an unsupervised learning task that aims to identify the samples deviating from common patterns. Typical algorithms include kernel density estimation, one-class SVM (scholkopf1999support), Isolation Forest (liu2008isolation), Deep SVDD (ruff2018deep), etc. Note that most of the dimensionality reduction methods (van2009dimensionality), such as PCA (abdi2010principal), KPCA (scholkopf1998nonlinear), and autoencoder (hinton2006reducing), can be used for anomaly detection or outlier detection. More papers about anomaly detection, especially those based on deep learning, can be found in (pang2021deep).
dietterich2018anomaly evaluated the detection performance of a few anomaly detection methods combined with different data imputation techniques. The experimental results showed that implementations contribute positively to the detection performance of unsupervised anomaly detection methods on incomplete data. fantii2022 investigated statistical process monitoring in the presence of missing data and introduced a fast incremental nonlinear matrix completion technique for online and sequential imputation. xu2021conformal developed ECAD, a method tailored for incomplete spatio-temporal data like traffic flow sequences, while ma2021probabilistic utilized Probabilistic Principal Component Analysis (PPCA) to recover missing values and identify anomalies. sarda2023unsupervised conducted a comparative analysis of existing unsupervised anomaly detection techniques applied to GAN-imputed datasets. These approaches follow a two-stage framework, in which the imputation models are trained exclusively on normal data or datasets with very few unlabeled outliers. Consequently, when applied to abnormal data during inference, the imputation model tends to reconstruct missing values based on learned normal patterns, thereby making anomalies appear more normal and reducing detection accuracy.
NEURIPS2024_f99f7b22 proposed an end-to-end method called ImAD for unsupervised anomaly detection in the presence of missing values. It integrates data imputation with anomaly detection into a unified optimization objective and automatically generates pseudo-abnormal samples to alleviate the imputation bias. They also proved the effectiveness of ImAD theoretically. zhang2025learning proposed an anomaly detection method for incomplete time series that flags unusual patterns by checking how consistently a two-step, patch-based imputation network fills the same series under different random masks.
@@ 7.4 Other Tasks on Incomplete Data
(sanguinetti2006missing; nguyen2008robust; hu2010semiparametric)
(de2018nonlinear; gilbert2018unsupervised; ling2021dimension; yan2024inference)
@@
(xie2012change; londschien2021change; zhao2022online; follain2022high)
(liu2025high; xu2025online; enikeeva2025change)
@@
(kallus2018causal; yang2019causal; pmlr-v89-tu19a; nguyen2020bayesian)
(wang2020causal; huang2020causal; hillis2021causal)
(gao2022missdag; levis2025robust; landsiedel2025causal)
reinforcement learning
(yamaguchi2020model; mei2023reinforcement; chasalow2025missing)
@@ 8.1 Effective Parameter Estimation
dempster1977maximum proved that the EM algorithm for handling missing data converges to a local maximum point of the likelihood function. Although this does not directly guarantee accurate missing data imputation, it guarantees the effectiveness of the EM algorithm in estimating the parameters of many models such as mixture models and robust regression. chandrasekher2020imputation studied the problem of single imputation for high-dimensional sparse linear regression and obtained optimal consistency rates of both the LASSO (ranstam2018lasso) and the square-root LASSO without modification. bertsimas2024simple studied the performance of impute-then-regress pipelines by contrasting theoretical and empirical evidence. It established the asymptotic consistency of such pipelines for a broad family of imputation methods and showed that mean-impute is asymptotically optimal. verchand2024high presented an asymptotically exact characterization of the risk of ridge-regularized logistic regression when the observed covariates stem from a Gaussian error-in-variables model.